The acceleration due to gravity on the Moon's surf
Answer #1

As, acceleration due to gravity = GM/R2

=>   acceleration due to gravity on moon = 16GM'/R2

Thus, for M' = M/96

=> Thus, mass of moon = M/96

where, M = mass of Earth